9. Given this information: • Beta: 1.6 • Market risk premium 8% • Risk free rate: 3% • Dividends expected to grow 4% per year • Last dividend: 2 EUR • Equity selling at 17.5 a) What is the expected cost of equity using the CAPM?

b) What is the expected cost of equity using the dividend growth model?

c) Is there any difference between them? And if risk free rate decrease by 1%?
Answer #1
Under the Capital Asset pricing model
Rs = Rf + Beta*(Rm-Rf)
Rs is the expected return on the security
where Rf is the risk free rate is 3%, Rm - Rf = difference between the expected return on the market
portfolio and the riskfree rate. (Rm-Rf)= 8%
Beta = 1.6
Rs = .03 + 1.6*(.08)
Rs = .158
a) The expected cost of equity using CAPM is 15.8%
D0 2
g2 0.04
D1 = D0*(1+g2)
D1 2.08
According to the dividend growth model.
P = D1/(R-g2)
where R is the expected cost of equity
P 17.5
17.5 = 2.08/(R-.04)
(R-.04) = 2.08/17.5
R - .04 = .11886
R = .11886 + .04
R = .15886
b) The expected cost of equity using the dividend growth model is
equal to 15.886%
c) There is no difference between the cost of equity using the CAPM and
the dividend growth model when the risk free rate is 3%.
When the risk free rate decreases by 1%, the Rf = 2%
Rm - Rf = 9%.
According to the CAPM,
Rs = .02+ 1.6*(.09)
Rs = 16.4%
The expected cost of equity using CAPM is 16.4%
The expected cost of equity using the dividend growth model is
equal to 15.886%
Difference = .164 - .15886
Difference = .00514.
Difference = .514%